OpenStudy (anonymous):

Please answer: 1. Determine whether a triangle can be formed with the given side lengths. If so, use Heron's formula to find the area of the triangle. a = 240 b = 132 c = 330 answers: No triangle is formed. 30,776.84 13,385.87 30,790.94 2. Express the complex number in trigonometric form. -4i answers: 4(cos 180° + i sin 180°) 4(cos 270° + i sin 270°) 4(cos 90° + i sin 90°) 4(cos 0° + i sin 0°) 3. Find a ⋅ b. a = 6i - 6j, b = 4i + 2j answers: 12 -36 <24, -12> <10, -4>

OpenStudy (anonymous):

which one you want to do?

OpenStudy (anonymous):

start with any please

OpenStudy (anonymous):

k lets do the first one

OpenStudy (anonymous):

you can definitely form a triangle with this, since the sum of any two sides is larger than the third now we can find the area

OpenStudy (anonymous):

first we need \(s=\frac{a+b+c}{2}\) in your case you get \[s=\frac{240+ 132+ 330 }{2}\] and a calculator

OpenStudy (anonymous):

i get \(s=351\) not done yet though

OpenStudy (anonymous):

now we get \[A=\sqrt{s(s-a)(s-b)(s-c)}\] which in your case is \[\sqrt{351(351-24)(351-132)(351-330)}\]

OpenStudy (anonymous):

think i lost @lanybug96

OpenStudy (anonymous):

im here! just putting together what you did

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

this is heron's formula \(A=\sqrt{s(s-a)(s-b)(s-c)}\) where \(s=\frac{a+b+c}{2}\)

OpenStudy (anonymous):

i would use a calculator, and i get this http://www.wolframalpha.com/input/?i= \sqrt{351%28351-240%29%28351-132%29%28351-330%29}

OpenStudy (anonymous):

made a typo there, put 24 where it should have been 240\[\sqrt{351(351-240)(351-132)(351-330)}\]

OpenStudy (anonymous):

maybe this link works better http://www.wolframalpha.com/input/?i=sqrt%28351%28351-240%29%28351-132%29%28351-330%29%29

OpenStudy (anonymous):

go with 13,385.87

OpenStudy (anonymous):

let me know if you want to do the next one too

OpenStudy (anonymous):

yeah okay that's what I got! :)

OpenStudy (anonymous):

good

OpenStudy (anonymous):



OpenStudy (anonymous):

here the absolute value is pretty clearly 4, and in any case all your answers have a 4 out front all you need is the angle

OpenStudy (anonymous):

that angle i hope is obvious, it is \(270\) since you are working in degrees

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

that tells you \(4i\) can be written as \[ 4(\cos (270°) + i \sin (270°)) \]

OpenStudy (anonymous):

oops i meant \(-4i\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you good with the dot product?

OpenStudy (anonymous):

yeah the last one I need help on to plz

OpenStudy (anonymous):

it is \(6\times 4+(-6)\times 2\)

OpenStudy (anonymous):

i get \(12\)

OpenStudy (anonymous):

your a life saver, thank you

OpenStudy (anonymous):

yw hope the method is clear in case you have to take a test or something these problems are kind of all over the map i assume it is an on line class